Solve the following problems. Use the properties of the sampling distribution of

sample means.

1. If a population has a mean of 5.7, what is the mean of the sampling distribution

of its means?

2. If a population has a variance of 7.4, what is the variance of the sampling

distribution of its means? The sampling distribution has a sample size of 2

and all possible samples are drawn with replacements.

3. If a population has a standard deviation of 3.2, what is the standard deviation

of the sampling distribution of its means? The sampling distribution has a

sample size of 4 and all possible samples are drawn with replacements.

4. If a population with size of 4 has a variance of 6.8, what is the variance of the

sampling distribution of its means? The sampling distribution has a sample size

of 3 and all possible samples are drawn without replacements.




1
Expert's answer
2021-05-07T09:39:13-0400
  1. "\mu=5.7" . The population mean is equal to the mean of the sampling distribution of its means. of its means.
  2. "\sigma_x^2=\frac{\sigma^2}{n}" ="\frac{7.4}{2}" =3.7
  3. "\sigma_x=\frac{\sigma}{\sqrt n}" ="\frac{3.2}{\sqrt 4}" =1.6
  4. "\sigma_x^2=\frac{\sigma^2}{n}" ="\frac{6.8}{3}" =2.67